# Lightweight Baselines for Medical Abstract Classification: DistilBERT with Cross-Entropy as a Strong Default

### v1

@@ Abstract
Large language models work well for many NLP tasks, but they are hard to deploy in health settings with strict cost, latency, and privacy limits. We revisit a lightweight recipe for medical abstract classification and ask how far compact encoders can go under a controlled budget. Using the public medical_abstracts corpus, we fine-tune BERT-base and DistilBERT with three objectives—standard cross-entropy, class-weighted cross-entropy, and focal loss—keeping tokenizer, sequence length, optimizer, and schedule fixed. DistilBERT with plain cross-entropy gives the best balance on the test set while using far fewer parameters than BERT-base. We report accuracy, Macro-F1, and Weighted-F1, release the evaluation code, and include confusion analyses to make error patterns clear. Our results suggest a practical default: start with a compact encoder and cross-entropy, then add calibration and task-specific checks before moving to heavier models.
@@ I Introduction
Biomedical literature is growing so fast that manual tracking is no longer feasible. This motivates reliable automated systems to triage, classify, and summarize large text collections. Automatic classification of medical abstracts supports literature search and clinical decision support. Large language models perform well [1], but their compute and memory costs limit use in healthcare settings with budget, latency, and privacy constraints (e.g., HIPAA) [2]. We therefore focus on strong, efficient, and reproducible baselines using compact encoders such as DistilBERT [3].
@@
Why compact baselines? Despite the excitement around large models, many real-world healthcare pipelines operate under strict cost-to-serve and governance requirements (on-prem, air-gapped or VPC-constrained deployments; auditable updates; predictable inference latencies). In such settings, compact encoders that are easy to fine-tune and calibrate often deliver a better accuracy–efficiency trade-off than heavyweight alternatives, particularly when the target task is well-posed and labeled at scale (e.g., abstract-level categorization). Our study therefore focuses on strong, reproducible baselines with minimal moving parts (two encoders $\times$ three loss functions), emphasizing clarity and deployability.
@@
Connection to domain-adaptive pretraining. A substantial body of work shows that domain-adaptive pretraining improves biomedical and scientific text modeling: SciBERT tailors representations to scientific prose [4]; BioBERT and PubMedBERT leverage large-scale biomedical corpora [5, 6]; and ClinicalBERT adapts to clinical narratives [7]. These models typically outperform purely general-domain encoders on in-domain tasks. Our objective here is complementary: we ask how far compact, readily available encoders (BERT-base and DistilBERT) can go on a public medical-abstract benchmark under identical, budgeted fine-tuning, and whether common imbalance remedies (class weighting, focal loss) are actually beneficial in this regime. Establishing a clean baseline is also useful for future comparisons against domain-specialized encoders.
Dataset characteristics and imbalance considerations. Abstracts condense background, methods, and outcome statements into short passages, often mixing rhetorical zones within sentences. The resulting label boundaries can be fuzzy even when class priors are only moderately skewed. In such circumstances, over-emphasizing ambiguous examples via reweighting or focal modulation may amplify noise and harm macro-F1, whereas plain cross-entropy (CE) can remain a surprisingly strong default. Our experiments revisit this trade-off in a controlled setting, reporting Accuracy, Macro-F1, and Weighted-F1, alongside confusion analyses and per-class profiles to make error structure explicit.
Calibration and governance. For decision-support applications, calibrated confidence matters as much as top-line accuracy. Prior studies show that pre-trained Transformers may exhibit miscalibration that is nonetheless amenable to simple post-hoc remedies (e.g., temperature scaling) without retraining [8]. We therefore discuss calibration metrics (ECE/Brier) and lightweight post-hoc adjustments in tandem with accuracy, aligning modeling practice with auditability and clinical governance [2].
Research questions. We ask: (RQ1) Among cross-entropy (CE), class-weighted CE (WCE), and focal loss (FL), which objective yields the best accuracy and macro-F1? (RQ2) Does a compact encoder (DistilBERT) match or surpass a heavier encoder (BERT-base) under identical budgets? (RQ3) How do objectives influence error structure across classes? (RQ4) How sensitive are baselines to practical knobs (sequence length, learning rate, early stopping)? (RQ5) What efficiency gains (parameters, on-disk size, throughput) accompany the best configuration?
@@ III-B Metrics
We report Accuracy, Macro-F1, and Weighted-F1 using scikit-learn implementations [21, 22]. In addition, we compute class-wise precision/recall/F1 and visualize error structure with confusion matrices and per-class bar charts [23, 24]. Unless otherwise noted, metrics are computed with the same evaluation code across all six configurations on the held-out test split.
@@ III-D Training and Implementation Details
For each of the six configurations (two encoders $\times$ three objectives), we export per-run JSON/CSV metrics (including global and per-class scores), confusion matrices, per-class metric plots, and model/tokenizer checkpoints. These artifacts enable exact reproduction of all numbers reported in Section IV.
@@ IV-A Experimental Setup
We evaluate six configurations formed by two encoders (BERT-base and DistilBERT) crossed with three objectives: cross-entropy (CE), class-weighted CE (WCE), and focal loss (FL). Unless noted, we fine-tune with AdamW (learning rate $2{\times}10^{-5}$), batch size $16$, for $3$ epochs; tokenization is uncased with truncation/padding at $256$ tokens. We adopt the official train/test split of medical_abstracts and reserve $10\%$ of the training set as a stratified validation split. All random seeds are fixed; we save the best checkpoint by validation Macro-F1 and evaluate on the test set. Metrics and plots are computed with scikit-learn utilities [21, 22, 23, 24]. The code relies on Transformers/Datasets [27, 28].
@@ IV-C Aggregate Performance
Table II reports the full scoreboard. DistilBERT with standard CE achieves the best balance (Accuracy $64.61\%$, Macro-F1 $64.38\%$, Weighted-F1 $63.25\%$), narrowly surpassing BERT-base with CE. Figure 1 visualizes the Macro-F1 ranking across all six settings. Figure 2 contrasts parameter count against Macro-F1 and shows that DistilBERT—the compact encoder distilled from BERT [3, 25]—occupies a favorable accuracy–efficiency region.
Observation O1 (Objective choice). For both encoders, WCE and FL underperform CE. The relative drops in Macro-F1 (vs CE) indicate that emphasizing difficult/minority examples does not translate into better global balance on this corpus, consistent with a balanced-to-moderately-imbalanced regime [18]. This aligns with broader evidence that when skew is mild and label ambiguity is present, cost-sensitive reweighting or focusing terms can over-amplify noisy/ambiguous instances and hurt precision [30, 11].
@@ IV-D Per-class Behavior and Error Structure
The two-column confusion panel (Figure 3) reveals strong diagonals for Class_1 and Class_4, while Class_5 shows distributed off-diagonal mass. The per-class bars (Figure 4) expose how objectives shift precision/recall:
@@ IV-G Takeaways for Practitioners
TABLE II: Overall results on medical_abstracts: Accuracy, Macro-F1, and Weighted-F1 (%).
Model Objective Acc. Macro-F1 Wtd-F1 bert-base-uncased cross_entropy 64.51 63.85 62.12 bert-base-uncased class_weight 62.88 62.43 59.66 distilbert-base-uncased cross_entropy 64.61 64.38 63.25 distilbert-base-uncased class_weight 62.29 62.22 59.24
Figure 1: Macro-F1 across six configurations; DistilBERT+CE leads.
Figure 2: Parameter count (M) vs Macro-F1. DistilBERT offers a favorable accuracy–efficiency trade-off.
Figure 3: Confusion matrices for all six configurations (top: BERT-base; bottom: DistilBERT).
@@ V Discussion, Limitations, and Impact
Cross-entropy performs better than class-weighted cross-entropy and focal loss because, under moderate class skew and label ambiguity, upweighting hard or minority examples can amplify noise and reduce precision; this matches findings that reweighting or resampling may shift rather than reduce confusions in mildly imbalanced text classification [30, 34, 11]. Focal loss helps in dense detection or long-tail vision tasks, but its $(1-p_{t})^{\gamma}$ factor can overemphasize uncertain points when ambiguity—not rarity—is the main issue [10]. While alternatives such as effective-number class-balanced loss and logit adjustment are reasonable in other regimes [32, 35], the label distribution of medical_abstracts and our fixed training budget make plain cross-entropy the most stable choice here [18].
@@ VI Conclusion and Future Work
We show that a simple, compute-efficient baseline, DistilBERT with cross-entropy, matches or slightly exceeds BERT-base at lower cost. On the medical_abstracts corpus [18, 19], class-weighting and focal loss do not help under the same budget. A practical path is to begin with a compact encoder and cross-entropy, then add calibration and explicit checks for dataset shift and governance before moving to heavier models [37, 38].
Future work includes extending to multimodal inputs from figures and tables while building strong safety and bias audits, informed by lightweight vision–language detectors and targeted bias stress suites [13, 14]. It is also useful to test compact encoders against larger systems reviewed in recent surveys and to examine scale versus clinical utility [39, 15], alongside efficient learning methods that lower parameters and cost [40]. To improve the workflow, agentic low-code frameworks can standardize pipelines and increase reproducibility [16], and topology-preserving projections can enrich error analysis beyond confusion matrices [17]. Architectural and optimization changes such as Bayesian-optimized attentive networks are another direction [41]. Finally, moving beyond static abstracts to longitudinal prediction [42] and exploring federated learning with data harmonization for privacy and non-IID settings remain important [43].

### v2

@@ Abstract
The research evaluates lightweight medical abstract classification methods to establish their maximum performance capabilities under financial budget restrictions. On the public medical_abstracts corpus, we fine-tune BERT-base and DistilBERT with three objectives—cross-entropy (CE), class-weighted CE, and focal loss—under identical tokenization, sequence length, optimizer, and schedule. DistilBERT with plain CE gives the strongest raw (argmax) trade-off, while a post-hoc operating-point selection (validation-calibrated, class-wise thresholds) substantially improves deployed performance; under this tuned regime, focal benefits most. We report Accuracy, Macro-F1, and Weighted-F1, release evaluation artifacts, and include confusion analyses to clarify error structure. The practical takeaway is to start with a compact encoder and CE, then add lightweight calibration/thresholding when deployment requires higher macro balance.
@@ I Introduction
Biomedical literature is expanding faster than manual curation can keep up, motivating automated systems for triage, classification, and summarization. While large language models perform well [1], their compute, latency, and privacy costs hinder deployment in healthcare settings governed by policies such as HIPAA [2]. We therefore emphasize strong, efficient, and reproducible baselines built on compact encoders such as DistilBERT [3].
@@
Why compact baselines? Despite the excitement around large models, many real-world healthcare pipelines operate under strict cost-to-serve and governance requirements (on-prem, air-gapped or VPC-constrained deployments; auditable updates; predictable inference latencies). In such settings, compact encoders that are easy to fine-tune and calibrate often deliver a better accuracy–efficiency trade-off than heavyweight alternatives, particularly when the target task is well-posed and labeled at scale (e.g., abstract-level categorization). Our study therefore focuses on strong, reproducible baselines with minimal moving parts (two encoders $\times$ three losses, plus a lightweight post-hoc operating-point selection) to maximize clarity and reproducibility.
@@
Relation to domain-adaptive pretraining. Domain-adaptive models typically improve in-domain text modeling: SciBERT targets scientific prose [4]; BioBERT and PubMedBERT leverage large biomedical corpora [5, 6]; ClinicalBERT adapts to clinical narratives [7]. Our question is complementary: how far can readily available compact encoders (BERT-base, DistilBERT) go on a public medical-abstract benchmark under identical, budgeted fine-tuning, and are common imbalance remedies (class weighting, focal loss) beneficial in this regime?
Dataset characteristics and imbalance. Abstracts mix background, methods, and outcomes, yielding fuzzy label boundaries even under moderate class skew. In such settings, reweighting or focal modulation may overemphasize ambiguous examples and harm macro balance; thus we treat plain cross-entropy (CE) as a strong raw default and empirically test alternatives.
Raw vs. deployed operating points. We report both raw (argmax) results and tuned results obtained via post-hoc operating-point selection: temperature scaling on validation [8] followed by class-wise threshold tuning, with thresholds frozen on test. Under raw decoding, CE is generally most stable; under tuned decoding, focal shows the largest gains—highlighting that training objectives and deployment thresholds should be evaluated separately.
Calibration and governance. Calibrated confidence is important for decision support; simple post-hoc calibration can mitigate miscalibration in pretrained Transformers without retraining [8], aligning modeling practice with auditability and privacy requirements [2].
Research questions. We study: (RQ1) which objective (CE, WCE, focal) best optimizes accuracy and Macro-F1; (RQ2) whether DistilBERT matches or surpasses BERT-base under identical budgets; (RQ3) how objectives shape per-class errors; (RQ4) sensitivity to practical knobs; (RQ5) efficiency gains (parameters, size, throughput); and (RQ6) which configuration performs best under a tuned deployment operating point.
@@ III-B Metrics
We report Accuracy, Macro-F1, and Weighted-F1 using scikit-learn implementations [21, 22]. In addition, we compute class-wise precision/recall/F1 and visualize error structure with confusion matrices and per-class bar charts [23, 24]. Unless otherwise noted, metrics are computed with the same evaluation code across all nine configurations on the held-out test split.
@@ III-D Training and Implementation Details
For each raw configuration (BERT/DistilBERT $\times$ {CE, WCE, FL}) and each tuned DistilBERT configuration (CE/WCE/FL), we export per-run JSON/CSV metrics (including global and per-class scores), confusion matrices, per-class metric plots, and model/tokenizer checkpoints, together with the selected thresholds for the tuned runs. These artifacts enable exact reproduction of all numbers reported in Section IV.
Operating Point Optimization (post-hoc)
After training, we optionally fit temperature scaling on the validation split [29] and perform a small grid search of class-wise decision thresholds to maximize validation Macro-F1. The selected temperature (often T=1T{=}1 when unhelpful) and thresholds are frozen and applied once to the test set (no refitting). We therefore report both Raw (argmax) and Tuned (class-wise thresholds, with temperature applied only when beneficial) results to decouple training quality from deployment-time operating-point choice.
@@ IV-A Experimental Setup
We evaluate six raw configurations formed by two encoders (BERT-base and DistilBERT) crossed with three objectives: cross-entropy (CE), class-weighted CE (WCE), and focal loss (FL). In addition, we report three tuned operating points for DistilBERT (CE/WCE/FL) obtained by post-hoc class-wise thresholding selected on validation and frozen on test. Unless noted, we fine-tune with AdamW (learning rate $2{\times}10^{-5}$), batch size $16$, for $3$ epochs; tokenization is uncased with truncation/padding at $256$ tokens. We adopt the official train/test split of medical_abstracts and reserve $10\%$ of the training set as a stratified validation split. All random seeds are fixed; we save the best checkpoint by validation Macro-F1 and evaluate on the test set. Metrics and plots are computed with scikit-learn utilities [21, 22, 23, 24]. The code relies on Transformers/Datasets [27, 28].
@@ IV-C Aggregate Performance
Table II summarizes results. Raw/argmax: DistilBERT with standard CE achieves the best balance (Accuracy $64.61\%$, Macro-F1 $64.38\%$, Weighted-F1 $63.25\%$), narrowly surpassing BERT-base with CE. Tuned (val-selected thresholds, frozen on test): focal loss benefits most from operating-point selection and attains the highest deployed scores (Acc. $\geq$ 80%), whereas CE also improves but to a lesser degree. (see Obs. O4) Figure 1 visualizes the Macro-F1 ranking across all nine settings. Figure 2 contrasts parameter count against Macro-F1 and shows that DistilBERT—the compact encoder distilled from BERT [3, 25]—occupies a favorable accuracy–efficiency region.
Observation O1 (Objective choice). Under raw/argmax decoding, WCE and FL underperform CE on this corpus, consistent with balanced-to-moderately-imbalanced label priors [18]. This aligns with broader evidence that when skew is mild and label ambiguity is present, cost-sensitive reweighting or focusing terms can over-amplify noisy/ambiguous instances and hurt precision [31, 11].
@@
Observation O4 (Decode mode matters). Objective rankings differ between raw and tuned settings: CE is the most stable raw default, while focal gains the most from post-hoc thresholds and yields the best deployed performance.
@@ IV-D Per-class Behavior and Error Structure
The $\times3$ confusion panel (Figure 3) reveals strong diagonals for Class_1 and Class_4, while Class_5 shows distributed off-diagonal mass. The per-class bars (Figure 4) expose how objectives shift precision/recall:
@@ IV-G Takeaways for Practitioners
TABLE II: Overall results on medical_abstracts: Accuracy, Macro-F1, and Weighted-F1 (%). We report raw (argmax) and tuned (validation-selected class-wise thresholds, frozen on test) results.
Configuration Acc. Macro-F1 Wtd-F1 Raw baselines (argmax) bert-base-uncased cross_entropy 64.51 63.85 62.12 bert-base-uncased class_weight 62.88 62.43 59.66 bert-base-uncased focal 62.81 62.47 59.93 distilbert-base-uncased cross_entropy 64.61 64.38 63.25 distilbert-base-uncased class_weight 62.29 62.22 59.24 distilbert-base-uncased focal 62.64 62.40 59.74 Post-hoc threshold-tuned (thresholds selected on val, fixed on test, optimized parameter 10e, 512, 3e-5, bs16) distilbert-base-uncased CE + tuned (all thresholds = 0.7) 74.03 70.73 72.42 distilbert-base-uncased Focal + tuned (all thresholds = 0.7) 82.22 77.55 81.22 distilbert-base-uncased ClassWeight + tuned (all thresholds = 0.7) 68.60 66.06 62.97
Figure 1: Macro-F1 for all settings. Left (gray): six raw configurations (BERT/DistilBERT $\times$ {CE, WCE, FL}). Right (blue): tuned configurations with class-wise thresholding (selected on validation and frozen on test). Tuned DistilBERT+Focal attains the best Macro-F1 (77.55%), tuned CE reaches 70.73%. See Table II for exact values.
Figure 2: Parameter count (millions) vs. Macro-F1. Color encodes loss (CE, Focal, Class Weight); marker shape distinguishes regimes (Raw, Tuned via per-class thresholding on val); fill indicates model (hollow=BERT $\approx$110M, solid=DistilBERT $\approx$66M). Tuned points lift Macro-F1 without changing parameter count.
Figure 3: Nine confusion matrices in a $\times3$ grid: top—BERT, middle—DistilBERT, bottom—DistilBERT (tuned). Columns: CE / WCE / FL.
@@ V Discussion, Limitations, and Impact
We decouple the training objective from the deployment operating point. Under moderate class skew and label ambiguity, upweighting hard/minority examples can amplify noise and reduce precision; accordingly, with raw/argmax decoding that isolates loss effects, cross-entropy (CE) remains the most stable and strongest on this corpus [31, 34, 11]. Focal loss is useful in long-tail vision settings [10], but its $(1-p_{t})^{\gamma}$ term may over-emphasize uncertain points when ambiguity—not rarity—is dominant. While alternatives like class-balanced losses or logit adjustment can help in different regimes [33, 35], the label distribution of medical_abstracts and our fixed budget make CE the most reliable raw choice [18]. Why then does focal score highest after tuning? The large gains arise from post-hoc operating-point selection (class-wise thresholds selected on validation and frozen on test), a generic step that can lift multiple models; this reflects improved decision thresholding rather than necessarily better pre-softmax representations. Given CE’s hyperparameter-free simplicity, robustness, and auditability, we retain it as the strong raw default, and recommend lightweight calibration/thresholding at deployment time when Macro-F1 is prioritized.
@@ VI Conclusion and Future Work
We show that a simple, compute-efficient baseline, DistilBERT with cross-entropy, matches or slightly exceeds BERT-base at lower cost. On the medical_abstracts corpus [18, 19], class-weighting and focal loss do not surpass CE under identical budgets with raw/argmax decoding; with post-hoc operating-point selection, focal yields the highest deployed scores, highlighting the practical value of threshold selection. A practical path is to begin with a compact encoder and cross-entropy, then add calibration and explicit checks for dataset shift and governance before moving to heavier models [37, 38].
Future work includes extending to multimodal inputs from figures and tables while building strong safety and bias audits, informed by lightweight vision–language detectors and targeted bias stress suites [13, 14]. It is also useful to test compact encoders against larger systems reviewed in recent surveys and to examine scale versus clinical utility [39, 15], alongside efficient learning methods that lower parameters and cost [40]. To improve the workflow, agentic low-code frameworks can standardize pipelines and increase reproducibility [16], and topology-preserving projections can enrich error analysis beyond confusion matrices [17]. Architectural and optimization changes such as Bayesian-optimized attentive networks are another direction [41]. Finally, moving beyond static abstracts to longitudinal prediction [42] and exploring federated learning with data harmonization for privacy and non-IID settings remain important [43]. A systematic study of operating-point selection across objectives and datasets—beyond temperature and simple threshold sweeps—remains an important direction for robust deployment.